The measure of one acute angle in a right triangle is four times the measure of the other acute angle.write and solve a system of equations to find the measure of the acute angles
Answer by fractalier(6550)

Call the first one x. The other must be 90-x, since they are in a right triangle. Then we have
x = 4(90-x)
x = 360 - 4x
5x = 360
x = 72 degrees and then
90-x = 18 degrees